Perpetual Inventory Using Weighted Average Beginning inventory, purchases, and sales for Meta-B1 are as follows: July 1 Inventory 100 units at $400 12 Sale 70 units 23 Purchase 120 units at $450 26 Sale 110 units a. Assuming a perpetual inventory system and using the weighted average method, determine the weighted average unit cost after the July 23 purchase. $ per unit b. Assuming a perpetual inventory system and using the weighted average method, determine the cost of the merchandise sold on July 26. $ c. Assuming a perpetual inventory system and using the weighted average method, determine the inventory on July 31. $

Perpetual Inventory Using Weighted Average

Beginning inventory, purchases, and sales for Meta-B1 are as follows:

July 1 Inventory 100 units at $400
12 Sale 70 units
23 Purchase 120 units at $450
26 Sale 110 units

a. Assuming a perpetual inventory system and using the weighted average method, determine the weighted average unit cost after the July 23 purchase.
$per unit

b. Assuming a perpetual inventory system and using the weighted average method, determine the cost of the merchandise sold on July 26.
$

c. Assuming a perpetual inventory system and using the weighted average method, determine the inventory on July 31.
$
Answer #1

a) Weighted average unit cost = (30*400+120*450)/(30+120) = 440 per unit

b) Cost of merchandise sold on july 26 = 440*110 = 48400

c) Ending inventory = 40*440 = 17600